# Picasso: Memory-Efficient Graph Coloring Using Palettes With Applications in Quantum Computing

### 12 Jan 2024 v1

@@ VIII Conclusions and Future Work
Our future work will focus on developing distributed multi-GPU parallel implementations along with new algorithms for predicting $\mathcal{P}$ and $\alpha$ values to enable better trade-offs for quality and performance. We plan to further optimize our algorithm to address inputs from diverse applications with varying degrees of sparsity. We also plan to develop efficient algorithms for clique partition that explore applications beyond quantum computing. To the best of our knowledge, this is the first of its kind study for computing coloring of dense graphs on limited memory accelerator platforms and believe that it will enable several applications that critically depend on the computation of clique partition and graph coloring, as well as enable the development of memory-efficient graph algorithms.

### 12 Feb 2024 v2

@@ VIII Conclusions and Future Work
Our future work will focus on developing distributed multi-GPU parallel implementations along with new algorithms for predicting $\mathcal{P}$ and $\alpha$ values to enable better trade-offs for quality and performance. We plan to further optimize our algorithm to address inputs from diverse applications with varying degrees of sparsity. We also plan to develop efficient algorithms for clique partitioning that explore applications beyond quantum computing. To the best of our knowledge, this is the first of its kind study for computing coloring of dense graphs on limited memory accelerator platforms and believe that it will enable several applications that critically depend on the computation of clique partitioning and graph coloring, as well as enable the development of memory-efficient graph algorithms.